In a random sample of eleven cell​ phones, the mean full retail price was ​$515.50 and the standard deviation was ​$212.00. Assume the population is normally distributed and use the​ t-distribution to find the margin of error and construct a 90​% confidence interval for the population mean mu. Interpret the results.
Answer #1

Solution :

t\alpha /2,df = 1.812

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 1.812 * (212 / \sqrt 11)

Margin of error = E = 115.8

The 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

515.5 - 115.8 < \mu < 515.5 + 115.8

399.7 < \mu < 631.3

A 90​% confidence interval for the population mean mu is : (399.7 , 631.3)